OpenStudy (anonymous):

What is the value of the function y= 2x-3 when x = –1? A. –5 B. 2 C. –1 D. 3

OpenStudy (aripotta):

just substitute -1 for x and solve for y

OpenStudy (anonymous):

y = 2x - 3......x = -1 y = 2(-1) - 3 y = -2-3 y = ?? can you finish ??